OpenStudy (anonymous):

when do we use quadratic formula? Can we use it whenever we want or there are special conditions when we use them?

jimthompson5910 (jim_thompson5910):

You can use the quadratic formula to solve any equation in the form \[\Large ax^2+bx+c=0\] where \[\Large a\neq 0\]

OpenStudy (zarkon):

It can also be used with functions like this... \[a(g(x))^2+b\cdot g(x)+c=0\]